# Controllable Flow Matching for Online Reinforcement Learning

### v1

@@ Introduction
Figure 2: The overall of CtrlFlow. The CFM model $\mathcal{M}$ is trained using data from environment replay buffer $\mathcal{B}_{\text{env}}$ (Left Top). During sampling (Left Bottom), $\mathcal{M}$ generates trajectory data while being guided by both the value guidance vector field $\mathcal{G}$ and control vector field $\mathcal{V}_{c}$. These components work in concert to refine the original vector field through an ODE-based process from $t=0$ to $t=1$. The resulting trajectories are then stored in the model replay buffer $\mathcal{B}_{\text{mod}}$, which is subsequently combined with $\mathcal{B}_{\text{env}}$ for policy learning (Right SAC).
@@
$\frac{\partial}{\partial t}\Phi^{t_{0},t}(\widehat{\tau}^{0})=v(\Phi^{t_{0},t}(\widehat{\tau}^{0})),\quad\Phi^{0}(\widehat{\tau}^{0})=\widehat{\tau}^{0}\in\mathbb{R}^{d},$ (6)
@@
Definition 6 (Non-linear Controllability Gramian Matrix)
@@
$\lambda_{\min}(N(u))\parallel\delta\widehat{\tau}\parallel^{2}\leq\delta\widehat{\tau}^{\top}N(u)\delta\widehat{\tau}\leq\lambda_{\max}(N(u))\parallel\delta\widehat{\tau}\parallel^{2},$ (10)
@@
Figure 4: Study on generated length and value guidance on MuJoCo Hopper-v3 task with five random seeds. (Left) Compare the performance of six different lengths with $h=2,5,8,10,30,50$ by model generation. (Right) The model with and without the value guidance vector field $\mathcal{G}$.
@@
Figure 5: (Top Left) Study on control theory on MoJoCo Hopper-v3 task. (Others) The comparison of controllable and cosine similarity between the generated final state and the ground truth.

### v2

@@ Introduction
Figure 2: The overall of CtrlFlow.
@@
$\frac{\partial}{\partial t}\Phi^{t_{0},t}(\widehat{\tau}^{0})=v(\Phi^{t_{0},t}(\widehat{\tau}^{0})),\quad\Phi^{0,0}(\widehat{\tau}^{0})=\widehat{\tau}^{0}\in\mathbb{R}^{d},$ (6)
@@
Definition 6 (Non-linear Controllability Gramian)
@@
$\lambda_{\min}(N(u))\parallel\delta\widehat{\tau}\parallel^{2}$ $\leq\delta\widehat{\tau}^{\top}N(u)\delta\widehat{\tau}$ (10)
$\leq\lambda_{\max}(N(u))\parallel\delta\widehat{\tau}\parallel^{2},$
@@
Figure 4: Study on generated length and value guidance on MuJoCo Hopper-v3 task with five random seeds in 50k steps. (Left) Compare the performance of six different lengths with $h=2,5,8,10,30,50$ by model generation. (Right) The model with and without the value guidance vector field $\mathcal{G}$.
@@
Figure 5: (Top Left) Study on control theory on MoJoCo Hopper-v3 task in 50k steps. (Others) The comparison of controllable and cosine similarity between the generated final state and the ground truth.